OpenStudy (anonymous):

5(8+h)^3-4(8+h)2+5=5h^3+116h^2+896h+2309 is the correct answer. i keep getting 5h^3+4h^2+56h+61. can someone tell me what im doing wrong?

OpenStudy (anonymous):

=5(512 + h^3 +24(8+h))-4(64 +h^2+16h)+5 =(2560 + 5h^3 +960h +120h^2) -(256 +4h^2 +64h) +5 =5h^3 + 116h^2 +896h +2309 formula- (a+b)^3= a^3+b^3+3ab(a+b) (a+b)^2=a^2 + b^2 +2ab